the weight of all male senior high school students from one province is measured and is found to be approximately normally distributed with a mean of 115 lbs and standard deviation of 3 lbs. find the percentile rank of the following weights?








A. 112 lbs








B. 115 lbs

1
Expert's answer
2022-03-09T06:43:06-0500

Population Mean "(\u03bc) =115"


Population Standard Deviation "(\u03c3) =3"


"Z=\\dfrac{X-\\mu}{\u03c3}\n\u200b"


(a) The percentile rank for 112 lbs


"Z=\dfrac{112-115}{3}"


"Z=-1.0"


From the standard normal distribution tables


"Z(-1.0)=0.15866"


=15.866th percentile


(b) The percentile rank for 115 lbs


"Z=\dfrac{115-115}{3}"


"Z=0"


From the standard normal distribution tables


"Z(0)=0.50000"


=50th percentile